# Linear Transformers are Versatile In-Context Learners

### v1

@@ Abstract
Recent research has demonstrated that transformers, particularly linear attention models, implicitly execute gradient-descent-like algorithms on data provided in-context during their forward inference step. However, their capability in handling more complex problems remains unexplored. In this paper, we prove that any linear transformer maintains an implicit linear model and can be interpreted as performing a variant of preconditioned gradient descent. We also investigate the use of linear transformers in a challenging scenario where the training data is corrupted with different levels of noise. Remarkably, we demonstrate that for this problem linear transformers discover an intricate and highly effective optimization algorithm, surpassing or matching in performance many reasonable baselines. We reverse-engineer this algorithm and show that it is a novel approach incorporating momentum and adaptive rescaling based on noise levels. Our findings show that even linear transformers possess the surprising ability to discover sophisticated optimization strategies.
Machine Learning, ICML
@@ 1 Introduction
Transformer architecture (Vaswani et al., 2017) has revolutionized the field of machine learning, driving breakthroughs across natural language processing, computer vision, and beyond. It has become the backbone of powerful foundation models (Anil et al., 2023; Achiam et al., 2023; Team et al., 2023; Jiang et al., 2023). However, despite their widespread success, the mechanisms that drive their performance remain an active area of research. A key component of their success is attributed to in-context learning (ICL, Brown et al., 2020) – an emergent ability of transformers to make predictions based on information provided within the input sequence itself, without explicit parameter updates.
@@
Specifically, von Oswald et al. (2023a) demonstrated that linear transformers can execute iterations of an algorithm similar to the gradient descent algorithm (which they call ${}^{++}$), with each attention layer representing one step of the algorithm. Later, Ahn et al. (2023); Zhang et al. (2023) further characterized this behavior, showing that the learned solution is a form of preconditioned GD, and this solution is optimal for one-layer linear transformers.
In this paper, we continue to study linear transformers trained on linear regression problems. We prove that any linear transformer maintains an implicit linear model. Under some restrictions, the algorithm it runs can be interpreted as a complex variant of preconditioned gradient descent with momentum-like behaviors.
While maintaining a linear model (regardless of the data) might seem restrictive, we nevertheless find that linear transformers can discover powerful optimization algorithms. As a first example, we prove that in case of ${}^{++}$, the preconditioner results in a second order optimization algorithm.
@@
Our experiments with two different noise variance distributions (uniform and categorical) demonstrate the remarkable flexibility of linear transformers. Training a linear transformer in these settings leads to an algorithm that outperforms ${}^{++}$ as well as various baselines derived from the exact closed-form solution of the ridge regression. We discover that this result holds even when training a linear transformer with diagonal weight matrices.
Through a detailed analysis, we reverse-engineered the learned algorithm, revealing key distinctions from ${}^{++}$, including momentum-like term and adaptive rescaling based on the noise levels.
@@ 2 Preliminaries
In this section we introduce notations for linear transformers, data, and type of problems that we consider.
@@ 2.1 Linear transformers and in-context learning
Given input sequence $e_{1},e_{2},...,e_{n}\in\mathbb{R}^{d}$, a single head in a linear self-attention layer is usually parameterized by 4 matrices, key $W_{K}$, query $W_{Q}$, value $W_{V}$ and projection $W_{P}$. The output of the non-causal layer at position $i$ is $e_{i}+\Delta e_{i}$ where $\Delta e_{i}$ is computed as
@@
If we have multiple heads $(P_{1},Q_{1}),(P_{2},Q_{2}),...,(P_{h},Q_{h})$, the effect is just the summation of all heads
@@
We define a linear transformer as a multi-layer neural network composed of $L$ linear self-attention layers parameterized by $\theta=\{Q^{l}_{k},P^{l}_{k}\}_{H,L}$. To isolate the core mechanisms, we consider a simplified decoder-only architecture, excluding MLPs and LayerNorm components. This architecture was also used in previous work (von Oswald et al., 2023a; Ahn et al., 2023).
@@
Inspired by von Oswald et al. (2023a), in this paper we consider a regression problem where data is provided as a sequence of tokens. Each token $e_{i}=(x_{i},y_{i})\in\mathbb{R}^{d+1}$ consists of a feature vector $x_{i}\in\mathbb{R}^{d}$ and its corresponding output $y_{i}\in\mathbb{R}$. Additionally, we append a query token $e_{n+1}=(x_{t},0)$ to the sequence, where $x_{t}\in\mathbb{R}^{d}$ represents test data. The goal of in-context learning is to predict $y_{t}$ for the test data $x_{t}$. We constrain the attention to only focus on the first $n$ tokens of the sequence so that it ignores the query token.
@@
$\Sigma=\sum_{i=1}^{n}x_{i}(x_{i})^{\top};$ $\alpha=\sum_{i=1}^{n}y_{i}x_{i};$ $\quad\lambda=\sum_{i=1}^{n}(y_{i})^{2}$
$\Sigma^{l}=\sum_{i=1}^{n}x^{l}_{i}(x^{l}_{i})^{\top};$ $\alpha^{l}=\sum_{i=1}^{n}y^{l}_{i}x^{l}_{i};$ $\quad\lambda^{l}=\sum_{i=1}^{n}(y^{l}_{i})^{2}$
@@ 2.2 Noisy regression model
where $e_{i}=(x_{i},y_{i})$ and $y_{i}=\langle w_{\tau},x_{i}\rangle+\xi_{i}$. This problem was initially explored by Garg et al. (2022). Later, von Oswald et al. (2023a) have demonstrated that a linear transformer (6) converges to a form of a gradient descent solution, which they called ${}^{++}$. We define this in details later.
@@
This scenario adds complexity because the model must predict $w_{\tau}$ for changing noise distribution, and the optimal solution likely would involve some sort of noise estimation.
Surprisingly, this simple modification leads to both Full and Diag linear transformers converging to a significantly better solution than ${}^{++}$. In fact, ${}^{++}$ fails to model noise variance estimation, instead converging to a solution which can be interpreted as a single noise variance estimate across all input data.
@@ 3 Related work
Our work is inspired by the line of work that demonstrated in-context learning can be viewed as (variants of) gradient descent (Akyürek et al., 2022; von Oswald et al., 2023a). For 1-layer linear transformer, several works (Zhang et al., 2023; Mahankali et al., 2023; Ahn et al., 2023) characterized the optimal parameters and training dynamics. More recent works extended the ideas to auto-regressive models (Li et al., 2023; von Oswald et al., 2023b) and nonlinear models (Cheng et al., 2023). Fu et al. (2023) noticed that transformers perform similarly to second-order Newton methods on linear data, for which we give a plausible explanation in Theorem 5.1.
@@
Other than the setting of linear models, several other works (Garg et al., 2022; Tarzanagh et al., 2023; Li et al., 2023; Huang et al., 2023; Tian et al., 2023a, b) considered optimization of transformers under different data and model assumptions. (Wen et al., 2023) showed that it can be difficult to interpret the “algorithm” performed by transformers without very strong restrictions.
@@
4 Linear transformers maintain linear models
While larger transformers with nonlinearities could represent complicated models, we show that linear transformers are restricted to maintaining a linear model based on the input, in the sense that the $l$-th layer output is always a linear function of the input with latent (and possibly nonlinear) coefficients.
@@ 4 Linear transformers maintain linear models
$x^{l+1}_{i}$ $=M^{l}x_{i}+y_{i}u^{l},$
$x^{l+1}_{t}$ $=M^{l}x_{t},$
$y^{l+1}_{i}$ $=a^{l}y_{i}-\langle w^{l},x_{i}\rangle,$
$y^{l+1}_{t}$ $=-\langle w^{l},x_{t}\rangle.$
This theorem implies that the output of linear transformer can always be explained as linear combinations of input with latent weights $a^{l}$ and $w^{l}$. This does not mean the matrices $M^{l}$, vectors $u^{l},w^{l}$ and numbers $a^{l}$ are linear. In fact they can be quite complex, which we characterize below:
@@
then one can recursively compute matrices $M^{l}$, vectors $u^{l},w^{l}$ and numbers $a^{l}$ for every layer with the following formula
@@
As we can see the updates to the parameters are complicated and nonlinear. This allows linear transformers to be creative and implement strong algorithms, as we will later see in Section 5. In fact, even when we restrict the $P$ and $Q$ matrices to be diagonal, linear transformers are still very flexible. The updates in this case can be further simplified to a more familiar form:
@@
In the setup of Theorem 4.1 with diagonal parameters (9), the parameters $u^{l},w^{l}$ are updated as
@@ 5 Power of diagonal attention matrices
In the previous section we saw that linear transformers are limited in the sense that they always maintain a linear model. However, this does not necessarily mean that they can’t solve some interesting in-context learning problems. Empirically, we have found that linear transformers are able to very accurately solve linear regression with mixed noise variance (7). Surprisingly, the final loss remains remarkably consistent even when the linear transformer’s $Q$ and $P$ matrices (3) are constrained to be diagonal. In this section we are going to study this special case and understand why they work so well.
@@
$\begin{split}w^{l}_{xx}&=\sum_{k=1}^{H}p^{l}_{x,k}q^{l}_{x,k},\quad w^{l}_{xy}=\sum_{k=1}^{H}p^{l}_{x,k}q^{l}_{y,k},\\ w^{l}_{yx}&=\sum_{k=1}^{H}p^{l}_{y,k}q^{l}_{x,k},\quad w^{l}_{yy}=\sum_{k=1}^{H}p^{l}_{y,k}q^{l}_{y,k}.\end{split}$ (9)
@@
$\begin{split}x^{l+1}_{i}&=x^{l}_{i}+w^{l}_{xx}\Sigma^{l}x^{l}_{i}+w^{l}_{xy}y^{l}_{i}\alpha^{l}\\ x^{l+1}_{t}&=x^{l}_{t}+w^{l}_{xx}\Sigma^{l}x^{l}_{t}+w^{l}_{xy}y^{l}_{t}\alpha^{l}\\ y^{l+1}_{i}&=y^{l}_{i}+w^{l}_{yx}\langle\alpha^{l},x^{l}_{i}\rangle+w^{l}_{yy}y^{l}_{i}\lambda^{l},\\ y^{l+1}_{t}&=y^{l}_{t}+w^{l}_{yx}\langle\alpha^{l},x^{l}_{t}\rangle+w^{l}_{yy}y^{l}_{t}\lambda^{l}.\end{split}$ (10)
In essence, four variables $w^{l}_{xx}$, $w^{l}_{xy}$, $w^{l}_{yx}$, $w^{l}_{yy}$ represent the flow of information between the data and the labels across layers. For instance, $w^{l}_{xx}$ measures how much information flows from $x^{l}$ to $x^{l+1}$, $w^{l}_{yx}$ measures the flow from $x^{l}$ to $y^{l+1}$ and so forth. Since the model can always be captured by these 4 variables, having many heads does not significantly increase its representation power. When there is only one head the equation $w^{l}_{xx}w^{l}_{yy}=w^{l}_{xy}w^{l}_{yx}$ is always true, while models with more than one head do not have this limitation. However empirically even models with one head is quite powerful.
5.1 ${}^{++}$ and least squares solver
${}^{++}$, introduced in von Oswald et al. (2023a), represents a linear transformer that is trained on a fixed noise variance problem (6). It is a variant of a diagonal linear transformer, with all the heads satisfying $q^{l}_{y,k}=0$. Dynamics are influenced only by $w^{l}_{xx}$ and $w^{l}_{yx}$, leading to simpler updates:
$\begin{split}x^{l+1}_{i}&=\left(I+w^{l}_{xx}\Sigma^{l}\right)x^{l}_{i}\\ y^{l+1}_{i}&=y^{l}_{i}+w^{l}_{yx}\langle\alpha^{l},x^{l}_{i}\rangle.\end{split}$ (11)
The update on $x$ acts as preconditioning, while the update on $y$ is just a gradient descent step based on the current data.
While existing analysis by Ahn et al. (2023) has not yielded fast convergence rates for ${}^{++}$, we show here that it is actually a second-order optimization algorithm for the least squares problem (4):

### v2

@@ Abstract
Recent research has demonstrated that transformers, particularly linear attention models, implicitly execute gradient-descent-like algorithms on data provided in-context during their forward inference step. However, their capability in handling more complex problems remains unexplored. In this paper, we prove that each layer of a linear transformer maintains a weight vector for an implicit linear regression problem and can be interpreted as performing a variant of preconditioned gradient descent. We also investigate the use of linear transformers in a challenging scenario where the training data is corrupted with different levels of noise. Remarkably, we demonstrate that for this problem linear transformers discover an intricate and highly effective optimization algorithm, surpassing or matching in performance many reasonable baselines. We analyze this algorithm and show that it is a novel approach incorporating momentum and adaptive rescaling based on noise levels. Our findings show that even linear transformers possess the surprising ability to discover sophisticated optimization strategies.
@@ 1 Introduction
The transformer architecture (Vaswani et al., 2017) has revolutionized the field of machine learning, driving breakthroughs across various domains and serving as a foundation for powerful models (Anil et al., 2023; Achiam et al., 2023; Team et al., 2023; Jiang et al., 2023). However, despite their widespread success, the mechanisms that drive their performance remain an active area of research. A key component of their success is attributed to in-context learning (ICL, Brown et al., 2020) – an emergent ability of transformers to make predictions based on information provided within the input sequence itself, without explicit parameter updates.
@@
Specifically, von Oswald et al. (2023a) demonstrated that linear transformers can execute iterations of an algorithm similar to the gradient descent algorithm (which they call GD++), with each attention layer representing one step of the algorithm. Later, Ahn et al. (2023); Zhang et al. (2023) further characterized this behavior, showing that the learned solution is a form of preconditioned GD, and this solution is optimal for one-layer linear transformers.
In this paper, we continue to study linear transformers trained on linear regression problems. We prove that each layer of every linear transformer maintains a weight vector for an underlying linear regression problem. Under some restrictions, the algorithm it runs can be interpreted as a complex variant of preconditioned gradient descent with momentum-like behaviors.
While maintaining a linear regression model (regardless of the data) might seem restrictive, we show that linear transformers can discover powerful optimization algorithms. As a first example, we prove that in case of GD++, the preconditioner results in a second order optimization algorithm.
@@
Our experiments with two different noise variance distributions (uniform and categorical) demonstrate the remarkable flexibility of linear transformers. Training a linear transformer in these settings leads to an algorithm that outperforms GD++ as well as various baselines derived from the exact closed-form solution of the ridge regression. We discover that this result holds even when training a linear transformer with diagonal weight matrices.
Through a detailed analysis, we reveal key distinctions from GD++, including momentum-like term and adaptive rescaling based on the noise levels.
@@ 2 Preliminaries
In this section we introduce notations for linear transformers, data, and type of problems we consider.
@@ 2.1 Linear transformers and in-context learning
Given input sequence $e_{1},e_{2},...,e_{n}\in\mathbb{R}^{d+1}$, a single head in a linear self-attention layer is usually parameterized by four matrices, key $W_{K}$, query $W_{Q}$, value $W_{V}$ and projection $W_{P}$. The output of the non-causal layer at position $i$ is $e_{i}+\Delta e_{i}$ where $\Delta e_{i}$ is computed as
@@
Multiple heads $(P_{1},Q_{1}),(P_{2},Q_{2}),...,(P_{h},Q_{h})$ simply sum their effects
@@
We define a linear transformer as a multi-layer neural network composed of $L$ linear self-attention layers parameterized by $\theta=\{Q^{l}_{k},P^{l}_{k}\}$ for $k=1\dots H,l=1\dots L$. To isolate the core mechanisms, we consider a simplified decoder-only architecture, excluding MLPs and LayerNorm components. This architecture was also used in previous work (von Oswald et al., 2023a; Ahn et al., 2023).
@@
Inspired by von Oswald et al. (2023a), in this paper we consider regression data as the token sequence. Each token $e_{i}=(x_{i},y_{i})\in\mathbb{R}^{d+1}$ consists of a feature vector $x_{i}\in\mathbb{R}^{d}$ and its corresponding output $y_{i}\in\mathbb{R}$. Additionally, we append a query token $e_{n+1}=(x_{t},0)$ to the sequence, where $x_{t}\in\mathbb{R}^{d}$ represents test data. The goal of in-context learning is to predict $y_{t}$ for the test data $x_{t}$. We constrain the attention to only focus on the first $n$ tokens of the sequence so that it ignores the query token.
@@
${\Sigma=\sum_{i=1}^{n}x_{i}(x_{i})^{\top};}$ $\quad{\alpha=\sum_{i=1}^{n}y_{i}x_{i};}$ ${\lambda=\sum_{i=1}^{n}(y_{i})^{2}}$
${\Sigma^{l}=\sum_{i=1}^{n}x^{l}_{i}(x^{l}_{i})^{\top};}$ $\quad{\alpha^{l}=\sum_{i=1}^{n}y^{l}_{i}x^{l}_{i};}$ ${\lambda^{l}=\sum_{i=1}^{n}(y^{l}_{i})^{2}}$
@@ 2.2 Noisy regression model
where $e_{i}=(x_{i},y_{i})$ and $y_{i}=\langle w_{\tau},x_{i}\rangle+\xi_{i}$. This problem was initially explored by Garg et al. (2022). Later, von Oswald et al. (2023a) have demonstrated that a linear transformer (6) converges to a form of a gradient descent solution, which they called GD++. We define this in details later.
@@
In other words, each training sequence $\tau$ has a fixed noise level $\sigma_{\tau}$, but different training sequences have different noise levels sampled from a specified distribution $p(\sigma_{\tau})$. This scenario adds complexity because the model must predict $w_{\tau}$ for changing noise distribution, and the optimal solution likely would involve some sort of noise estimation. We have found that empirically, GD++ fails to model this noise variance and instead converges to a solution which can be interpreted as a single noise variance estimate across all input data.
@@ 3 Related work
Our work builds on research that frames in-context learning as (variants of) gradient descent (Akyürek et al., 2022; von Oswald et al., 2023a). For 1-layer linear transformer, several works Zhang et al. (2023); Mahankali et al. (2023); Ahn et al. (2023) characterized the optimal parameters and training dynamics. More recent works extended the ideas to auto-regressive models (Li et al., 2023; von Oswald et al., 2023b) and nonlinear models (Cheng et al., 2023). Fu et al. (2023) noticed that transformers perform similarly to second-order Newton methods on linear data, for which we give a plausible explanation in Theorem 5.1.
@@
Other than the setting of linear models, several other works (Garg et al., 2022; Tarzanagh et al., 2023; Li et al., 2023; Huang et al., 2023; Tian et al., 2023a, b) considered optimization of transformers under different data and model assumptions. Wen et al. (2023) showed that it can be difficult to interpret the “algorithm” performed by transformers without very strong restrictions.
@@
4 Linear transformers maintain linear regression model at every layer
While large, nonlinear transformers can model complex relationship, we show that linear transformers are restricted to maintaining a linear regression model based on the input, in the sense that the $l$-th layer output is always a linear function of the input with latent (and possibly nonlinear) coefficients.
@@ 4 Linear transformers maintain linear regression model at every layer
$x^{l+1}_{i}$ $=M^{l}x_{i}+y_{i}u^{l},$ $x^{l+1}_{t}$ $=M^{l}x_{t},$
$y^{l+1}_{i}$ $=a^{l}y_{i}-\langle w^{l},x_{i}\rangle,$ $y^{l+1}_{t}$ $=-\langle w^{l},x_{t}\rangle.$
Note that $M^{l}$, $u^{l},w^{l}$ and $a^{l}$ are not linear in the input, but this still poses restrictions on what the linear transformers can do. For example we show that it cannot represent a quadratic function:
Theorem 4.2.
Suppose the input to a linear transformer is $(x_{1},y_{1}),(x_{2},y_{2}),...,(x_{n},y_{n})$ where $x_{i}\sim N(0,I)$ and $y_{i}=w^{\top}x_{i}$, let the $l$-th layer output be $(x^{l}_{1},y^{l}_{1}),(x^{l}_{2},y^{l}_{2}),...,(x^{l}_{n},y^{l}_{n})$ and let $y^{l}=(y^{l}_{1},...,y^{l}_{n})$ and $y^{*}=(x_{1}(1)^{2},x_{2}(1)^{2},...,x_{n}(1)^{2})$ (here $x_{i}(1)$ is just the first coordinate of $x_{i}$), then when $n\gg d$ with high probability the cosine similarity of $y^{*}$ and $y^{l}$ is at most 0.1.
Theorem 4.1 implies that the output of linear transformer can always be explained as linear combinations of input with latent weights $a^{l}$ and $w^{l}$. The matrices $M^{l}$, vectors $u^{l},w^{l}$ and numbers $a^{l}$ are not linear and can in fact be quite complex, which we characterize below:
@@
then one can recursively compute matrices $M^{l}$, vectors $u^{l},w^{l}$ and numbers $a^{l}$ for every layer using
@@
The updates to the parameters are complicated and nonlinear, allowing linear transformers to implement powerful algorithms, as we will later see in Section 5. In fact, even with diagonal $P$ and $Q$, they remain flexible. The updates in this case can be further simplified to a more familiar form:
@@
In the setup of Theorem 4.1 with diagonal parameters, $u^{l},w^{l}$ are updated as
@@ 5 Power of diagonal attention matrices
Although linear transformers are constrained, they can solve complex in-context learning problems. Empirically, we have found that they are able to very accurately solve linear regression with mixed noise variance (7), with final learned weights that are very diagonal heavy with some low-rank component (see Fig. 4). Surprisingly, the final loss remains remarkably consistent even when their $Q$ and $P$ matrices (3) are diagonal. Here we will analyze this special case and explain its effectiveness.
@@
$\begin{split}\omega^{l}_{xx}&=\textstyle{\sum_{k=1}^{H}p^{l}_{x,k}q^{l}_{x,k},}\quad\textstyle{\omega^{l}_{xy}=\sum_{k=1}^{H}p^{l}_{x,k}q^{l}_{y,k},}\\ \omega^{l}_{yx}&=\textstyle{\sum_{k=1}^{H}p^{l}_{y,k}q^{l}_{x,k},}\quad\textstyle{\omega^{l}_{yy}=\sum_{k=1}^{H}p^{l}_{y,k}q^{l}_{y,k}.}\end{split}$ (9)
@@
$\begin{split}x^{l+1}_{i}&=x^{l}_{i}+\omega^{l}_{xx}\Sigma^{l}x^{l}_{i}+w^{l}_{xy}y^{l}_{i}\alpha^{l}\quad\quad\quad\,\,\\ x^{l+1}_{t}&=x^{l}_{t}+\omega^{l}_{xx}\Sigma^{l}x^{l}_{t}+w^{l}_{xy}y^{l}_{t}\alpha^{l}\\ y^{l+1}_{i}&=y^{l}_{i}+\omega^{l}_{yx}\langle\alpha^{l},x^{l}_{i}\rangle+\omega^{l}_{yy}y^{l}_{i}\lambda^{l},\quad\quad\\ y^{l+1}_{t}&=y^{l}_{t}+\omega^{l}_{yx}\langle\alpha^{l},x^{l}_{t}\rangle+\omega^{l}_{yy}y^{l}_{t}\lambda^{l}.\end{split}$ (10)
Four variables $\omega^{l}_{xx}$, $\omega^{l}_{xy}$, $\omega^{l}_{yx}$, $\omega^{l}_{yy}$ represent information flow between the data and the labels across layers. For instance, the term controlled by $\omega^{l}_{xx}$ measures information flow from $x^{l}$ to $x^{l+1}$, $\omega^{l}_{yx}$ measures the flow from $x^{l}$ to $y^{l+1}$ and so forth. Since the model can always be captured by these 4 variables, having many heads does not significantly increase its representation power. When there is only one head the equation $\omega^{l}_{xx}\omega^{l}_{yy}=\omega^{l}_{xy}\omega^{l}_{yx}$ is always true, while models with more than one head do not have this limitation. However empirically even models with one head is quite powerful.
5.1 GD++ and least squares solver
GD++, introduced in von Oswald et al. (2023a), represents a linear transformer that is trained on a fixed noise variance problem (6). It is a variant of a diagonal linear transformer, with all the heads satisfying $q^{l}_{y,k}=0$. Dynamics are influenced only by $\omega^{l}_{xx}$ and $\omega^{l}_{yx}$, leading to simpler updates:
$\begin{split}x^{l+1}_{i}&=\left(I+\omega^{l}_{xx}\Sigma^{l}\right)x^{l}_{i}\\ y^{l+1}_{i}&=y^{l}_{i}+\omega^{l}_{yx}\langle\alpha^{l},x^{l}_{i}\rangle.\end{split}$ (11)
The update on $x$ acts as preconditioning and the update on $y$ performs gradient descent on the data.
While existing analysis by Ahn et al. (2023) has not yielded fast convergence rates for GD++, we show here that it is actually a second-order optimization algorithm for the least squares problem (4):